Operations Report – Nov 10th

Crew 214 Operations Report 10-11-2019
SOL: 0
Name of person filing report: Shane Usher

Non-nominal systems: Opportunity, Suit 10, Suit 5 Helmet, No power to multiple downstairs wall sockets.
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning, Suit 5 Helmet falling apart

Generator: run
Hours run: 12h
From what time last night: 1915
To what time this morning: 0729
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 82%

Diesel Reading – 55%
Station Propane Reading – 74%
Ethanol Free Gasoline: N/A

Water (loft tank): 23 gallons
Water Meter: 1457085 units
Water (static tank): 150 gallons (not filled)
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 109.6
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 87%
Currently charging: yes. Note that opening hood to clean mud from engine compartment loosened power cable, that has now been addressed. The power cable was connected to un-powered wall sockets and was an obstruction in hab rear entrance. Cable has been re-routed out of the way to power powerboard above Hab rear air-lock.

Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Curiosity rover used: yes
Hours: 115.2
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? Still in Bicknel for service
CrewCar used and why, where?. CrewCar: 9 Nov MDRS Odometer 179069 miles, 9 Nov Refuelled at Hollow Mountain, 9 Nov picked up Crew 215 at Grand Junction 179241, 9 Nov refueled at Hollow Mountain, 9 Nov returned to MDRS, 10 Nov Refuelled at Hollow Mountain, 10 Nov to Grand Junction for Crew 214

General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Suit 2 has poor airflow, Suit 5 Helmet falling apart, Suit 10 is not functioning will investigate and seek guidance from CAPCOM.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (16L). The heater automatically came on at 5 pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: Shane Usher – mild headache, staying hydrated, 2 panadol

Questions, concerns and requests to Mission Support: Static tank not filled. Water heater unusable. Opportunity is currently unusable.

Copyright © The Mars Society. All rights reserved. | Main Site